OpenStudy (anonymous):

solve x-1/4=3/7. please show work

OpenStudy (anonymous):

x = 3/7 +1/4 x = 19/28

OpenStudy (anonymous):

x-(1/4)=3/7 x+(1/4)-(1/4)=(3/7)+(1/4) x=(3/7)+(1/4) x=(3/7)(4/4)+(1/4)(7/7) x=(12/28)+(7/28) x=(12+7)/28 x=19/28